OpenStudy (anonymous):

if \(A \cap B\), \(A' \cap B\), \(B \cap A'\) are independent events, how to prove that \(A' \cap B'\) are also independent? I know that they are independent but I don't know how to prove.

OpenStudy (zarkon):

I think you need to clarify the problem. it doesn't look right

OpenStudy (anonymous):

one set is not "independent"

OpenStudy (anonymous):

What do we mean by independent here?

OpenStudy (anonymous):

you need two sets. A and B are independent means \[P(A|B)=P(A)\]

OpenStudy (anonymous):

or if you prefer \[\frac{P(A\cap B)}{P(B)}=P(A)\]

OpenStudy (zarkon):

it is probably best to write \[P(A\cap B)=P(A)P(B)\] just in case \(P(B)=0\)

OpenStudy (zarkon):

we do have this...if A and B are independent then A and B' are independent A' and B are independent A' and B' are independent here the 'and' is not the intersection

OpenStudy (anonymous):

yes i suppose. then again if P(B) = 0 then A and B are independent for all A

OpenStudy (anonymous):

i am going to bet that zarkon stated the question correctly.

OpenStudy (zarkon):

\[P(A'\cap B')\] \[=1-P(A\cup B)\] \[=1-(P(A)+P(B)-P(A\cap B))\] \[=1-P(A)-P(B)+P(A\cap B)\] \[=1-P(A)-P(B)+P(A)P(B)\] \[=1-P(A)-P(B)(1-P(A))\] \[=(1-P(A))(1-P(B))\] \[=P(A')P(B')\]

OpenStudy (anonymous):

yeah zarkon stated it correctly lol

OpenStudy (anonymous):

sorry for that